# On $f$-Divergence Principled Domain Adaptation: An Improved Framework

### v1

@@ D.10 Generalization Bounds based on $\chi^{2}$-DD
$\mathrm{D}^{h,\mathcal{H}_{r}}_{\chi^{2}}(\nu||\mu)=\sup_{h^{\prime},t}t\left({{\mathbb{E}_{\nu}}\left[g(X)\right]}-{{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)-\frac{t^{2}\mathrm{Var}_{\mu}g(X)}{4}.$
@@
${{\mathbb{E}_{\nu}}\left[g(X)\right]}-C^{\prime}{{\mathbb{E}_{\mu}}\left[g(X)\right]}\leq\inf_{t\geq 0}\frac{1}{t}\mathrm{D}^{h,\mathcal{H}_{r}}_{\chi^{2}}(\nu||\mu)+\frac{t\mathrm{Var}_{\mu}g(X)}{4}-C{{\mathbb{E}_{\mu}}\left[g(X)\right]}.$
When $t$ and $C$ satisfy $\frac{t\mathrm{Var}_{\mu}g(X)}{4}-C{{\mathbb{E}_{\mu}}\left[g(X)\right]}\leq 0$, we have ${{\mathbb{E}_{\nu}}\left[g(X)\right]}-{{\mathbb{E}_{\mu}}\left[g(X)\right]}\leq\inf_{t,C}\frac{1}{t}\mathrm{D}^{h,\mathcal{H}_{r}}_{\chi^{2}}(\nu||\mu)+C{{\mathbb{E}_{\mu}}\left[g(X)\right]}$.
@@
$\leq$ $\sup_{h^{\prime},t}t\left({{\mathbb{E}_{\nu}}\left[g(X)\right]}-{{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)-\frac{t^{2}\mathrm{Var}_{\mu}g(X)}{4}-t\left({{\mathbb{E}_{\hat{\nu}}}\left[g(X)\right]}-{{\mathbb{E}_{\hat{\mu}}}\left[g(X)\right]}\right)+\frac{t^{2}\mathrm{Var}_{\hat{\mu}}g(X)}{4}$
$=$ $\sup_{h^{\prime},t}t\left({{\mathbb{E}_{\nu}}\left[g(X)\right]}-{{\mathbb{E}_{\hat{\nu}}}\left[g(X)\right]}\right)+t\left({{\mathbb{E}_{\hat{\mu}}}\left[g(X)\right]}-{{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)+\frac{t^{2}}{4}\left(\mathrm{Var}_{\hat{\mu}}g(X)-\mathrm{Var}_{\mu}g(X)\right)$
@@
$\qquad+\sup_{h^{\prime}}\frac{t_{0}^{2}}{4}\left(\mathrm{Var}_{\hat{\mu}}g(X)-\mathrm{Var}_{\mu}g(X)\right).$
@@
$\sup_{h^{\prime}}\frac{t_{0}^{2}}{4}\left(\mathrm{Var}_{\hat{\mu}}g(X)-\mathrm{Var}_{\mu}g(X)\right)$
@@
D.11 Proof of Proposition 1
@@ Appendix E More Experiment Details and Additional Results
We adopt the experimental setup from Acuna et al. (2021) and build upon their publicly available code, accessible at https://github.com/nv-tlabs/fDAL/tree/main. Additionally, we leverage some settings from the implementations of Zhang et al. (2019) and Long et al. (2018), which can be found at https://github.com/thuml/MDD/tree/master and https://github.com/thuml/CDAN, respectively.
@@
Acuna et al. (2021) also investigate the empirical performance of combining $f$-DAL with a sampling-based implicit alignment technique from Jiang et al. (2020), enhancing their original $f$-DAL. However, our findings reveal that our $f$-DD consistently outperforms $f$-DAL with implicit alignment, as detailed in Table 6. This observation suggests that the performance gain can be achieved by simply adopting a more tightly defined variational representation. Furthermore, in Table 6, the entry labeled $f$-DD (Best) corresponds to the average of the maximum accuracy across KL-DD, $\chi^{2}$-DD, and Jeffereys-DD for each subtask. Notably, this aggregated result shows slight improvement on Office-31 when compared to the individual performance of Jeffereys-DD.
@@
As outlined in Section 6, instead of invoking a stochastic gradient-based optimizer (e.g., SGD) to update $t$, we utilize a quadratic approximation for the optimal $t$ as presented in Birrell et al. (2022). The approximately optimal KL-DD (OptKL-DD) is expressed as follows:
@@
This approximation is obtained by Birrell et al. (2022) through a Taylor expansion around $t=1$, with additional details provided in their Appendix B. In our implementation of OptKL-DD, similar to KL-DD, we use $\hat{\ell}$ to replace $\ell$.
@@
To visualize model representations (output of $h_{\rm rep}$) trained with $f$-DAL, $\chi^{2}$-DD, KL-DD, and Jeffereys-DD, we leverage t-SNE (Van der Maaten & Hinton, 2008). In Figure 3, we present visualization results using USPS (U) as the source domain and MNIST (M) as the target domain. Notably, while the original $f$-DAL already provides satisfactory results, our $f$-DD achieves further improvements in representation alignment.

### v2

@@ D.11 Generalization Bounds based on $\chi^{2}$-DD
$\mathrm{D}^{h,\mathcal{H}_{r}}_{\chi^{2}}(\nu||\mu)=\sup_{h^{\prime},t}t\left({{\mathbb{E}_{\nu}}\left[g(X)\right]}-{{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)-\frac{t^{2}\mathrm{Var}_{\mu}\left(g(X)\right)}{4}.$
@@
${{\mathbb{E}_{\nu}}\left[g(X)\right]}-C^{\prime}{{\mathbb{E}_{\mu}}\left[g(X)\right]}\leq\inf_{t\geq 0}\frac{1}{t}\mathrm{D}^{h,\mathcal{H}_{r}}_{\chi^{2}}(\nu||\mu)+\frac{t\mathrm{Var}_{\mu}\left(g(X)\right)}{4}-C{{\mathbb{E}_{\mu}}\left[g(X)\right]}.$
When $t$ and $C$ satisfy $\frac{t\mathrm{Var}_{\mu}\left(g(X)\right)}{4}-C{{\mathbb{E}_{\mu}}\left[g(X)\right]}\leq 0$, we have ${{\mathbb{E}_{\nu}}\left[g(X)\right]}-{{\mathbb{E}_{\mu}}\left[g(X)\right]}\leq\inf_{t,C}\frac{1}{t}\mathrm{D}^{h,\mathcal{H}_{r}}_{\chi^{2}}(\nu||\mu)+C{{\mathbb{E}_{\mu}}\left[g(X)\right]}$.
@@
$\leq$ $\sup_{h^{\prime},t}t\left({{\mathbb{E}_{\nu}}\left[g(X)\right]}-{{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)-\frac{t^{2}\mathrm{Var}_{\mu}\left(g(X)\right)}{4}-t\left({{\mathbb{E}_{\hat{\nu}}}\left[g(X)\right]}-{{\mathbb{E}_{\hat{\mu}}}\left[g(X)\right]}\right)+\frac{t^{2}\mathrm{Var}_{\hat{\mu}}\left(g(X)\right)}{4}$
$=$ $\sup_{h^{\prime},t}t\left({{\mathbb{E}_{\nu}}\left[g(X)\right]}-{{\mathbb{E}_{\hat{\nu}}}\left[g(X)\right]}\right)+t\left({{\mathbb{E}_{\hat{\mu}}}\left[g(X)\right]}-{{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)+\frac{t^{2}}{4}\left(\mathrm{Var}_{\hat{\mu}}\left(g(X)\right)-\mathrm{Var}_{\mu}\left(g(X)\right)\right)$
@@
$\qquad+\sup_{h^{\prime}}\frac{t_{0}^{2}}{4}\left(\mathrm{Var}_{\hat{\mu}}\left(g(X)\right)-\mathrm{Var}_{\mu}\left(g(X)\right)\right).$
@@
$\sup_{h^{\prime}}\frac{t_{0}^{2}}{4}\left(\mathrm{Var}_{\hat{\mu}}\left(g(X)\right)-\mathrm{Var}_{\mu}\left(g(X)\right)\right)$
@@
D.12 Threshold Learning Example
Consider a popular threshold learning example.
Example 1 (Threshold Learning).
Let $\mathcal{X}=\mathbb{R}$. A threshold function, denoted as $h_{c}$, is defined s.t. $h_{c}(x)=0$ if $x<c$ and $h_{c}(x)=1$ if $x\geq c$. Let the source domain $\mu$ be an uniform distribution on $[0,1]$ and let the target domain $\nu$ be another uniform distribution on $[0,2]$. Let =$\mathcal{H}=\{h_{c}|c\in[0,\frac{1}{2}]\}$ and assume the ground-truth hypothesis is $h^{*}_{\frac{1}{2}}$ for the both the source and target domains. Let $\ell$ be the zero-one loss.
Recall that $\mathrm{D}^{h,\mathcal{H}}_{\rm KL}(\nu||\mu)=\sup_{h^{\prime},t}t\mathbb{E}_{\nu}[\ell(h,h^{\prime})]-\log\mathbb{E}_{\mu}e^{t\ell(h,h^{\prime})}$. Hence,
$\mathrm{D}^{h_{\frac{1}{2}},\mathcal{H}}_{\rm KL}(\nu||\mu)=\sup_{t}\int_{0}^{\frac{1}{2}}\frac{t}{2}\cdot 1dx-\log\left(\int_{0}^{\frac{1}{2}}{e^{t}}dx+\int_{\frac{1}{2}}^{1}{e^{0}}dx\right)=\sup_{t}\frac{t}{4}-\log\left(\frac{1}{2}+\frac{e^{t}}{2}\right)=0.131.$
Now let $r=1/4$, so $\mathcal{H}_{r}=\{h_{c}|c\in[\frac{1}{4},\frac{1}{2}]\}$, by a similar calculation we have
$\mathrm{D}^{h_{\frac{1}{2}},\mathcal{H}_{\frac{1}{4}}}_{\rm KL}(\nu||\mu)=\sup_{t}\frac{t}{8}-\log(\frac{3}{4}+\frac{e^{t}}{4})=0.048\text{ and }R^{r}_{\mu}=\frac{1}{4}.$
Based on Remark 5.3, we can set $C_{2}=0.1$ and $C_{1}=3.74$. Hence, we have
$0.27\cdot\mathrm{D}^{h_{\frac{1}{2}},\mathcal{H}_{\frac{1}{4}}}_{\rm KL}(\nu||\mu)+0.1\cdot R^{r}_{\mu}=0.038\leq\sqrt{\mathrm{D}^{h_{\frac{1}{2}},\mathcal{H}}_{\rm KL}(\nu||\mu)}=0.36$
Since $h^{*}_{\frac{1}{2}}\in\mathcal{H}_{r}\subset\mathcal{H}$, we have $\lambda^{*}=\lambda^{*}_{r}=0$ in this case. This example justifies that the localization technique can significantly tighten the bound. Note that $C_{2}=0.1$ and $C_{1}=3.74$ are not the optimal choice so localized bound can be even tighter.
In addition, as a more extreme case, let $r\to 0$ so $\mathcal{H}_{r}=\{h_{\frac{1}{2}}\}$. Then by a similar calculation, we have $\mathrm{D}^{h_{\frac{1}{2}},\mathcal{H}_{0}}_{\rm KL}(\nu||\mu)=R^{r}_{\mu}=0$, while $\mathrm{D}^{h_{\frac{1}{2}},\mathcal{H}}_{\rm KL}(\nu||\mu)$ remains unchanged. Notice that $R_{\nu}(h_{\frac{1}{2}})-R_{\mu}(h_{\frac{1}{2}})=0$ so localized $f$-DD gives a tightest target error bound in this case.
D.13 Proof of Proposition 1
@@ Appendix E More Experiment Details and Additional Results
We adopt the experimental setup from [1] and build upon their publicly available code, accessible at https://github.com/nv-tlabs/fDAL/tree/main. Additionally, we leverage some settings from the implementations of [9] and [49], which can be found at https://github.com/thuml/MDD/tree/master and https://github.com/thuml/CDAN, respectively. Our code is available at https://github.com/ZiqiaoWangGeothe/f-DD.
Figure 2: Illustration of the adversarial training framework for $f$-DD-based UDA. The framework includes the representation network ($h_{\rm rep}$), the main classifier ($h_{\rm cls}$), and the auxiliary classification network ($h^{\prime}_{\rm cls}$). It jointly minimizes the empirical risk on the source domain and the approximated $f$-DD between the source and target domains.
@@
[1] also investigate the empirical performance of combining $f$-DAL with a sampling-based implicit alignment technique from [51], enhancing their original $f$-DAL. However, our findings reveal that our $f$-DD consistently outperforms $f$-DAL with implicit alignment, as detailed in Table 6. This observation suggests that the performance gain can be achieved by simply adopting a more tightly defined variational representation. Furthermore, in Table 6, the entry labeled $f$-DD (Best) corresponds to the average of the maximum accuracy across KL-DD, $\chi^{2}$-DD, and Jeffereys-DD for each subtask. Notably, this aggregated result shows slight improvement on Office-31 when compared to the individual performance of Jeffereys-DD.
@@
As outlined in Section 6, instead of invoking a stochastic gradient-based optimizer (e.g., SGD) to update $t$, we utilize a quadratic approximation for the optimal $t$ as presented in [23]. The approximately optimal KL-DD (OptKL-DD) is expressed as follows:
@@
This approximation is obtained by [23] through a Taylor expansion around $t=1$, with additional details provided in their Appendix B. In our implementation of OptKL-DD, similar to KL-DD, we use $\hat{\ell}$ to replace $\ell$.
@@
To visualize model representations (output of $h_{\rm rep}$) trained with $f$-DAL, $\chi^{2}$-DD, KL-DD, and Jeffereys-DD, we leverage t-SNE [52]. In Figure 4, we present visualization results using USPS (U) as the source domain and MNIST (M) as the target domain. Notably, while the original $f$-DAL already provides satisfactory results, our $f$-DD achieves further improvements in representation alignment.
More Advance Network Architecture
While our empirical study primarily aims to compare with [1] and improve their algorithms, in this section, we also evaluate our $f$-DD on more complicated models, which can serve as a valuable reference for practitioners.
In particular, we update our original backbone, ResNet-50, with pretrained transformer-based models. Specifically, we use the pretrained Vision transformer (ViT) base model (vit-base-patch16-224) [75] and pretrained Swin-based transformer (swin-base-patch4-window7-224) [76], while keeping all other settings in our algorithms unchanged. The results on Office-31 are presented in Table 10.
Table 10: Accuracy (%) on the Office-31 benchmark.
ViT-based [75] 91.2 99.2 100.0 90.4 81.1 80.6 91.1
SSRT-ViT [77] 97.7 99.2 100.0 98.6 83.5 82.2 93.5
PMTrans-ViT [78] 99.1 99.6 100.0 99.4 85.7 86.3 95.0
Ours-ViT 98.0 $\pm$ 0.2 99.2 $\pm$ 0.0 100 $\pm$ 0.0 98.6 $\pm$ 0.2 83.5 $\pm$ 0.5 83.9 $\pm$ 0.4 93.9
Swin-based [76] 97.0 99.2 100.0 95.8 82.4 81.8 92.7
PMTrans-Swin [78] 99.5 99.4 100.0 99.8 86.7 86.5 95.3
Ours-Swim 98.7 $\pm$ 0.2 99.2 $\pm$ 0.0 100 $\pm$ 0.0 99.2 $\pm$ 0.2 86.1 $\pm$ 0.1 85.6 $\pm$ 0.3 94.8
Results on models other than ours are taken directly from [78, Table 2]. Here ViT-based and Swin-based refer to source-only training using ViT and Swin transformers, respectively, and “Ours” particularly refers to our weighted-Jeffereys discrepancy. While our method does not outperform the current SOTA, PMTrans, (especially for ViT), but close to it, this should demonstrate the practical utility of the proposed approach. Notably, our approach not only significantly improves the respective backbones but also outperforms the strongest baseline, SSRT-ViT [77], as compared in [78].
The fact the our approach does not outperform PMTrans should come at no surprise. In particular, PMTrans and all other compared models all include while our approach not only neglected these ingredients but also simply use hyperparameter based on our earlier ResNet-50 settings without further tuning.
One notable difference between our method and other SOTA approaches is the absence of any pseudo-labeling strategy in our framework. Specifically, our theoretical framework is constructed under the assumption of zero knowledge of target labels. In a strategy that successfully exploits pseudo-labelling, usually explicit or implicit prior knowledge are available (e.g., for the purpose of selecting hyperparameters relating to pseudo labeling). For example, such knowledge may present itself as a labeled validation set in the target domain. Such knowledge, if available, may change our problem setup and require a refinement of our theory. We anticipate that combining our method with advanced pseudo-labeling strategies, data augmentation techniques like Mixup (which itself inherently relies on pseudo-labels), and label smoothing will improve our results, possibly approaching the current SOTA performance.